Use the following data to answer the below question.

Required reserve ratio is​ 10%

Checkable deposits $900

Savings Deposits $181

Total reserves $188

Currency held by the public $139

What is the level of required​ reserves? $90

What is the level of excess​ reserves? $98

What is the value of the monetary base ​$ 327

What is the currency ​ratio? (round to 2 decimals) 0.15

What is the money​ multiplier? (round to 2 decimals)
Answer #1

Level of required​ reserves=the required reserve ratio times the amount of checkable deposits

=10% * $900= $90

Level of required​ reserves=$90

The level of excess​ reserves = The total amount of reserves - required​ reserves

=$188 -$90 =$98

The level of excess​ reserves =$98

The value of the monetary base =Currency ( C) + reserves ( R )

=$139 +$188 = $327

The value of the monetary base =$327

The currency ​ratio = The desired level of currency C / checkable deposits D

=$139/$900 =0.15

The currency ​ratio =0.15

The money​ multiplier=

m = (1 + c) / (r + e + c)

r = Required reserve ratio = 10% =0.10

c= The currency ​ratio=0.15

e = {excess​ reserves / checkable deposits} =  $98/$900 =0.11

The money​ multiplier= (1 + c) / (r + e + c)

=(1+0.15) / (0.10+0.11+0.15)

=1.15/1.36 = 0.85

The money​ multiplier= 0.85